OpenStudy (anonymous):

Choose the equation that could be used to find two consecutive integers whose sum is 67. Answer n + (n + 1) = 67 n - (n + 1) = 67 n + (n + 2) = 67 n + n = 67 Choose the equation that could be used to find two consecutive integers whose sum is 67. Answer n + (n + 1) = 67 n - (n + 1) = 67 n + (n + 2) = 67 n + n = 67 @Mathematics

OpenStudy (diyadiya):

n + (n + 1) = 67

OpenStudy (amistre64):

ida used 67/2

OpenStudy (anonymous):

first one

OpenStudy (amistre64):

"that could be used" ... first and last seem to be on the board to me

OpenStudy (amistre64):

33+33 = 66 34+34 = 68 seems to me that it could then be 33 and 34 to average the 2 lol

OpenStudy (anonymous):

\[ n + (n + 1) = 67 \Rightarrow 2n+1=67 \Rightarrow n=33 \] so the two consecutive numbers are 33 and 34.

OpenStudy (anonymous):

first is right because n+n+1=67 on solving gives 33 so 33 and 34